# Reinforcing Multimodal Understanding and Generation with Dual Self-rewards

### v2

@@ 5 Conclusion
In this work, we proposed a dual self-reward mechanism to enhance the generation and understanding capabilities of LMMs in an efficient, self-supervised, and unified manner. By leveraging the dual self-reward mechanism, we designed a unified optimization strategy that jointly optimizes the two capabilities of the unified LMM, leading to substantial improvements in both generation and understanding tasks. We also investigated the performance of different optimization strategies, including the joint optimization strategy with a unified model and the adversarial-like optimization strategy with two separate models. Extensive experiments on various visual understanding and generation benchmarks demonstrate the effectiveness of our proposed method. Overall, our method showcases the potential of improving the generation and understanding capabilities of LMMs in a self-supervised and unified manner, paving the way for future research on multimodal alignment.
Limitations
@@ Appendix A Technical Appendices
A.1 Qualitative Comparison
We present qualitative comparisons of Janus-Pro-7B and Janus-Pro-7B tuned with DSR on visual generation and understanding tasks in Figure 8 and Figure 9, respectively. The images and responses are sampled from the models with the same setup, where the temperature is set to 1.0 and the random seed is set to 42.
@@ A.1 Qualitative Comparison
Figure 8: Qualitative comparison of visual generation. We sample multiple images for same text prompts using Janus-Pro-7B and reinforced Janus-Pro-7B with the proposed DSR.
Figure 9: Qualitative comparison of visual understanding. We sample multiple responses for same image with a question using Janus-Pro-7B and reinforced Janus-Pro-7B with the proposed DSR.
A.2 Demonstration of DSR
@@
Figure 10: Demonstration of our dual self-reward for visual generation. We sample multiple images for a given text prompt and then compute DSR for the images. In this demonstration, the first generated image is the best one, which is also the one with the highest DSR.
Figure 11: Demonstration of our dual self-reward for visual understanding. We sample multiple text descriptions for a given image and then compute DSR for the descriptions. In this demonstration, the upper generated description is better, which is also the one with the higher DSR.
A.3 Experiments with Show-o
We conduct additional experiments with another backbone, Show-o [5], to further validate the effectiveness of our method.
Show-o is a unified multimodal model that unifies autoregressive and discrete diffusion modeling for inputs and outputs in different modalities. It applies full attention for image tokens and causal attention for text tokens, and it is trained with a combination of next token prediction (NTP) and mask token prediction (MTP) objectives. To ensure efficiency, we follow diffu-GRPO [62], using one-step unmasking with randomly droped input text tokens condition to obtain the estimated log-probability of the generated image tokens.
@@ A.3 Experiments with Show-o
Table 5: Evaluation on T2I-CompBench. Und. and Gen. denote “understanding” and “generation”. Higher ($\uparrow$) values indicate better performance. The best score is in bold, with the second best score underlined. Line in blue is our tuned model. The line with is the results we reproduce for fair comparison.
@@
StrucDiffusion [47] 0.4990 0.4218 0.4900 0.1386 0.3111 0.3355
CompDiffusion [48] 0.4063 0.3299 0.3645 0.0800 0.2980 0.2898
Attend&Excite [49] 0.6400 0.4517 0.5963 0.1455 0.3109 0.3401
PixArt-$\alpha$ [50] 0.6690 0.4927 0.6477 0.2064 0.3197 0.3433
CoMat [41] 0.7827 0.5329 0.6468 0.2428 0.3187 0.3680
SD-v1.5 [14] 0.3758 0.3713 0.4186 0.1165 0.3112 0.3047
SD-XL-base-1.0 [51] 0.5879 0.4687 0.5299 0.2131 0.3119 0.3237
FLUX.1 [52] 0.7407 0.5718 0.6922 0.2863 0.3127 0.3703
@@
Show-o [5] 0.56 0.41 0.46 0.20 0.30 0.29
Show-o [5] 0.5892 0.4002 0.4489 0.2268 0.3026 0.3058
@@
A.4 Experiments using CLIP as Reward Model
We conduct additional experiments using CLIP [63] as the reward model, following SimpleAR [64]. We only optimize the visual generaton capability of Janus-Pro-7B with SimPO, and the results are shown in Table 6 and Table 7.
Table 6: Evaluation on T2I-CompBench. Und. and Gen. denote “understanding” and “generation”, respectively. Higher ($\uparrow$) values indicate better performance. Line in blue is our tuned model. The line with is the results reproduced by us for fair comparison.
[…]
Janus-Pro-7B [7] 0.6426 0.3487 0.4848 0.2061 0.3086 0.3510
Janus-Pro-7B + CLIP 0.7700 0.5609 0.7320 0.2302 0.3144 0.3863
[…]
Table 7: Evaluation on understanding benchmarks.
[…]
Janus-Pro-7B [7] 37.0 74.0 86.8 79.6 71.9 41.1
Janus-Pro-7B + CLIP 36.9 74.0 86.6 79.6 71.9 41.3
Janus-Pro-7B + DSR 38.9 81.5 86.6 80.1 71.9 41.1
A.5 Details of the Understanding Benchmarks
@@
MMMU [46]
includes multi-choice questions and open-ended QA, with 62 and 988 cases, respectively.

### v3

@@ 5. Conclusion
In this work, we proposed a self-improving framework SUDER, with a dual self-reward mechanism to enhance the generation and understanding capabilities of LMMs in an efficient, self-supervised, and unified manner. By leveraging the dual self-reward mechanism, we designed a unified optimization strategy that jointly optimizes the two capabilities of the unified LMM, leading to substantial improvements in both generation and understanding tasks. We also investigated the performance of different optimization strategies, including the joint optimization strategy with a unified model and the adversarial-like optimization strategy with two separate models. Extensive experiments on various visual understanding and generation benchmarks demonstrate the effectiveness of our proposed method. Overall, our method showcases the potential of improving the generation and understanding capabilities of LMMs in a self-supervised and unified manner, paving the way for future research on multimodal alignment.
6. Limitations
@@ Appendix A Technical Appendices
A.1. Implementation Details
A.2. Qualitative Comparison
We present qualitative comparisons of Janus-Pro-7B and Janus-Pro-7B tuned with SUDER on visual generation and understanding tasks in Figure 7 and Figure 8, respectively. The images and responses are sampled from the models with the same setup, where the temperature is set to 1.0 and the random seed is set to 42.
@@ A.2. Qualitative Comparison
Figure 7. Qualitative comparison of visual generation. We sample multiple images for the same text prompts using Janus-Pro-7B and reinforced Janus-Pro-7B with the proposed SUDER.
Figure 8. Qualitative comparison of visual understanding. We sample multiple responses for same image with a question using Janus-Pro-7B and reinforced Janus-Pro-7B with the proposed DSR.
A.3. Demonstration of DSR
@@
Figure 9. Demonstration of our dual self-reward for visual generation. We sample multiple images for a given text prompt and then compute DSR for the images. In this demonstration, the first generated image is the best one, which is also the one with the highest DSR.
Figure 10. Demonstration of our dual self-reward for visual understanding. We sample multiple text descriptions for a given image and then compute DSR for the descriptions. In this demonstration, the upper generated description is better, which is also the one with the higher DSR.
A.4. Experiments with Show-o
We conduct additional experiments with another backbone, Show-o (Xie et al., 2024), to further validate the effectiveness of our method.
Show-o is a unified multimodal model that unifies autoregressive and discrete diffusion modeling for inputs and outputs in different modalities. It applies full attention for image tokens and causal attention for text tokens, and it is trained with a combination of next token prediction (NTP) and mask token prediction (MTP) objectives. To ensure efficiency, we follow diffu-GRPO (Zhao et al., 2025), using one-step unmasking with randomly droped input text tokens condition to obtain the estimated log-probability of the generated image tokens.
@@ A.4. Experiments with Show-o
Table 6. Evaluation on T2I-CompBench. Und. and Gen. denote “understanding” and “generation”. Higher ($\uparrow$) values indicate better performance. The best score is in bold, with the second best score underlined. Line in blue is our tuned model. The line with is the results we reproduce for fair comparison.
@@
StrucDiffusion (Feng et al., 2022) 0.4990 0.4218 0.4900 0.1386 0.3111 0.3355
CompDiffusion (Liu et al., 2022) 0.4063 0.3299 0.3645 0.0800 0.2980 0.2898
Attend&Excite (Chefer et al., 2023) 0.6400 0.4517 0.5963 0.1455 0.3109 0.3401
PixArt-$\alpha$ (Chen et al., 2024a) 0.6690 0.4927 0.6477 0.2064 0.3197 0.3433
CoMat (Jiang et al., 2024) 0.7827 0.5329 0.6468 0.2428 0.3187 0.3680
SD-v1.5 (Rombach et al., 2022) 0.3758 0.3713 0.4186 0.1165 0.3112 0.3047
SD-XL-base-1.0 (Podell et al., 2023) 0.5879 0.4687 0.5299 0.2131 0.3119 0.3237
FLUX.1 (Labs, 2024) 0.7407 0.5718 0.6922 0.2863 0.3127 0.3703
@@
Janus-Pro-7B (Chen et al., 2025) 0.6426 0.3487 0.4848 0.2061 0.3086 0.3510
[…]
Show-o (Xie et al., 2024) 0.56 0.41 0.46 0.20 0.30 0.29
Show-o (Xie et al., 2024) 0.5892 0.4002 0.4489 0.2268 0.3026 0.3058
@@
A.5. Details of the Understanding Benchmarks